# Evolutionary Policy Optimization

### v2

@@ Abstract
Figure 1: We introduce Evolutionary Policy Optimization (EPO) (Left), a novel policy gradient algorithm that integrates a genetic algorithm (GA) with policy gradients. During training, a population of agents, each represented by a latent embedding (gene) and sharing a common actor-critic network, interacts with the environment (Middle). Among them, one agent is designated as the master agent, which is trained using amortized experience aggregated from all agents in the population to enhance learning efficiency and stability (Right).
@@ 1 Introduction
Reinforcement learning (RL) has become a powerful paradigm for training autonomous decision-making agents across a wide range of domains, including games [40, 2, 47, 30], robotics [19, 1, 22, 17], and large-language-model alignment [33, 13, 46]. Like all trial-and-error methods, RL is inherently sample-inefficient, and the problem is particularly acute for on-policy algorithms, which discard past experience after each update.
@@
Despite this drawback, on-policy model-free methods—commonly referred to as policy gradients [38, 30]—are widely adopted in real-world applications [40, 2, 19, 1, 33]. Their popularity stems from their ability to achieve higher asymptotic returns and stable performance, making them the method of choice in data-rich settings such as simulations and game engines [26], where large batches are easy to generate. Unfortunately, recent works [41, 34] show that policy-gradient algorithms do not scale well with larger batch sizes: because data are collected from the current policy, adding more parallel environments does not guarantee greater diversity. Instead, the data distribution quickly converges, so simply running more environments on additional GPUs yields highly correlated trajectories rather than useful variety.
@@
Evolutionary algorithms (EAs) contribute the complementary strength of explicit diversity through random perturbations, making them natural partners for RL. Their combination—evolutionary RL (EvoRL) [31, 36]—maintains a population of policies and refines them with evolutionary operators. Yet EvoRL has not been widely adopted. Classic EvoRL relies on gradient-free evolutionary strategies (ES) [20, 3, 48], whose simplest variants reduce to random search [27]: they parallelize well but remain extremely sample-inefficient. Methods such as CEM-RL [35] introduce off-policy updates to improve efficiency but degrade asymptotic performance owing to distribution shift. Population-Based Training (PBT) [34, 18] ensembles multiple on-policy agents, yet each policy evolves in isolation, wasting the opportunity to amortize experience across the population.
@@
Can we design an algorithm that (i) scales with the number of parallel environments, (ii) learns efficiently from shared experience, and (iii) preserves the high asymptotic returns and stability of on-policy methods? We answer in the affirmative. We introduce Evolutionary Policy Optimization (EPO), a novel policy-gradient algorithm that merges the strengths of EA and policy gradients: EA supplies diverse experience during massively parallel training, and experience from all agents is amortized to train a master agent via the split-and-aggregate policy-gradient (SAPG) formulation [41], which performs off-policy updates with importance sampling to fold follower data into the master’s on-policy set.
@@
EPO maintains a population of agents alongside a master agent, all sharing the same network weights (analogous to gene-expression rules). This design prevents parameter bloat as the population scales, enabling efficient training and storage of large networks. Each agent is conditioned on a unique latent variable (“gene”), ensuring behavioral diversity while preserving coherence. Periodically, Darwinian selection [8] is applied: low performers are removed, and elite agents are retained. These elites undergo crossover and mutation, injecting controlled diversity without excessive divergence. All agents are updated synchronously with policy gradients using shared reward signals. In every iteration, EPO aggregates experience from all agents into the master via SAPG updates [41], enabling the master to learn from diverse, high-quality data and fully exploit massive parallelism.
@@
We evaluate EPO on several challenging and realistic environments spanning manipulation (Isaac Gym Allegro-Kuka [26]), locomotion (ANYmal Walking [16], Unitree A1 Parkour [4, 49]), and classic control benchmarks (DeepMind Control Suite [45]). Across this suite, EPO significantly outperforms prior state-of-the-art RL methods. Moreover, our scaling-law analysis shows that EPO effectively scales with increased computational resources, making it a promising approach for large-scale RL training.
@@ 2 Related Work
On-policy approaches [43] update the policy using data collected by the current policy, ensuring that the training data remains closely aligned with the policy’s behavior. This strategy often leads to higher asymptotic rewards and stable performance, but at the cost of sample inefficiency, as past experiences from older policies are discarded. Despite this drawback, on-policy methods are the preferred choice in settings where data is abundant, such as simulations and game engines, where large training batches can be efficiently generated. In practice, on-policy RL has proven highly effective for training autonomous decision-making agents in games [40, 2, 47, 30], robotics [19, 1, 22, 17], and more recently, in developing reasoning capabilities in large language models (LLMs) [33, 13, 46]. Two widely used algorithms in this category are Trust Region Policy Optimization (TRPO) [37] and Proximal Policy Optimization (PPO) [38], both designed to stabilize learning by constraining the magnitude of policy updates.
@@
Unlike on-policy algorithms, off-policy methods allow for policy updates using data collected from different or older policies, enabling more efficient reuse of past experiences. Classic approaches such as Q-learning and its deep variant Deep Q-Network (DQN) [29] rely on a replay buffer to store state-action transitions for more stable learning. More recent continuous control algorithms, such as Deep Deterministic Policy Gradient (DDPG) [25, 39], Twin Delayed DDPG (TD3) [11], and Soft Actor-Critic (SAC) [14], extend these ideas within an actor-critic framework, further improving stability and performance. While off-policy methods are typically more sample-efficient than on-policy methods, they often suffer from distribution shift and overestimation of action values, which can hinder stable learning. To effectively amortize experience across all policies in a population while ensuring stable updates, we adopt the Split and Aggregate Policy Gradient (SAPG) formulation [41], which performs off-policy updates via importance sampling, enabling the aggregation of experience from individual follower policies into the on-policy data of a master policy.
@@
Another related line of work is Evolutionary Reinforcement Learning (EvoRL) [31], which integrates population-based search with policy optimization to explore a broad range of candidate solutions. Historically, evolutionary computation (EC) has been employed to optimize neural network weights [44, 5, 6], architectures [12], and hyperparameters [18, 42]. For instance, OpenAI ES [7] applied a streamlined evolution strategy to Atari games. More recently, researchers have integrated gradient-based updates alongside evolutionary operations (e.g., mutation and crossover), improving scalability and leveraging the representational power of deep neural networks [35, 21]. For example, [20] extends Deep Deterministic Policy Gradient (DDPG) by introducing evolutionary operations on a population of policies, combining the exploratory benefits of evolution with gradient-based optimization for more robust learning. However, these methods remain vulnerable to distribution shift, which undermines their asymptotic performance. In this work, we propose to combine the strengths of evolutionary search with on-policy updates, enabling higher asymptotic rewards while maintaining high sample efficiency.
@@ 3 Preliminaries
is one of the most well-known Evolutionary Algorithms (EAs). It relies on three key operators: selection, mutation, and crossover. A core concept in genetic algorithms is the population, where each individual represents a potential solution evaluated using a fitness function. The classic genetic algorithm emphasizes crossover as the primary mechanism for exploration [15] and is usually applied to deal with the problem of hyper-parameter tuning in reinforcement learning.
@@ 4 Approach
To further enhance learning efficiency, we adopted a hybrid policy update strategy [41]. The master agent updates its parameters using both (i) on-policy data collected in the most recent episode and (ii) off-policy data sampled from the trajectories of the entire population. Meanwhile, each non-master agent is updated using only its own on-policy data. This hybrid learning process ensures both efficient exploration and sample reuse, making EPO well-suited for large-scale reinforcement learning. In the following sections, we describe each component of EPO in detail.
@@ 4.1 Genetic Algorithm for Agent Selection
The design philosophy of our algorithm is twofold: (1) to generate diverse yet bounded data for training the master agent and (2) to ensure that non-master agents are updated efficiently, enabling them to produce useful yet diverse data. Drawing inspiration from biological evolution, we propose initializing a population of agents ($\{\pi_{k}\}_{k=1}^{K}$), where each agent is associated with a unique latent embedding ($\{\phi_{k}\}_{k=1}^{K}$). All agents share a common actor-critic network, parameterized by $\theta$ and $\psi$, respectively. The actor is conditioned on both the current state and its latent embedding to generate actions, while the value function is similarly conditioned to estimate value functions. Prior works have also leveraged latent-conditioned policies to encourage diverse behaviors [9, 10]. Among these agents, one is designated as the master agent—for simplicity, we define the master agent as $\pi_{1}(\theta,\psi,\phi_{1})$, though the general framework allows for alternative assignments.
@@ 4.2 Hybrid-Policy Optimization
Different agents generate diverse data through varied behaviors, which can be beneficial if the master agent can effectively learn from all available data. However, since data distributions collected by different agents may vary significantly, incorporating off-policy data requires careful correction. Following [28, 41], we apply importance sampling to reweight updates from different policies. Specifically, we sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ collected by agents $\{\pi_{k}\}_{k=2}^{K}$ to get $S^{\prime}_{1}$. To update the master agent $\pi_{1}$ using $S^{\prime}_{1}$, we define the off-policy loss as:
@@
Algorithm 1 Evolutionary Policy Optimization
2:  Initialize $K$ empty cyclic replay buffers $\{S\}_{i=1}^{K}$
3:  Initialize $N$ environments $\{E\}_{i=1}^{N}$
4:  for k = 1, …, $K$ do
5:     $E^{k}=\{E\}_{i=\frac{(k-1)N}{K}}^{\frac{kN}{K}}$triangleright Assign environments
6:  end for
7:  for iteration=1,2,… do
8:     for k = 2, …, $K$ do
9:        $f_{k}\leftarrow Evaluate(E^{k},\pi_{k})$
10:     end for
11:     if Eq 3 then
12:        Rank the population based on fitness scores $f_{k}$
13:        Select top $x$ agents as elites to form Set $X$, append to $Y$
14:        while $|Y|<K-1$ do
15:           ϕsuperscriptitalic-$\phi^{\prime}\leftarrow$Crossover between randomly sampled $\phi_{i},\phi_{j}$,where $\pi_{i}(\theta,\psi,\phi_{i}),\pi_{j}(\theta,\psi,\phi_{j})\in X$
16:           ϕ′′superscriptitalic-$\phi^{\prime\prime}\leftarrow$ Mutate($\phi^{\prime}$)
17:           Append $\pi(\theta,\psi,\phi^{\prime\prime})$ to $Y$
18:        end while
19:     end if
20:     $\{\pi_{i}\}_{i=2}^{K}\leftarrow Y$
21:     for k = 1, …, $K$ do
22:        CollectData$(E^{k},\pi_{k})$
23:     end for
24:     $L\leftarrow 0$
25:     Sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ to get $S_{1}^{\prime}$
26:     $L\leftarrow L+$ OffPolicyLoss($S_{1}^{\prime}$)
27:     for k = 1, …, $K$ do
28:        $L\leftarrow L+$ OnPolicyLoss($S_{k}$)
29:     end for
30:     Update $\theta$, $\psi$, $\{\phi_{k}\}_{k=1}^{K}$
31:  end for
32:  Return $\theta$, $\psi$, $\phi_{1}$
@@ 5 Experiments
Figure 2: We evaluate our algorithm on seven challenging environments that span a diverse set of tasks, including manipulation [34], locomotion [4], and classic control benchmarks [45]
We evaluate our method on eight challenging and realistic environments, including manipulation [34], locomotion [16, 4], and classic control benchmarks [45], comparing it against state-of-the-art (SOTA) approaches. Additionally, we conduct ablation studies to assess the contribution of each component. To demonstrate scalability, we analyze our method’s performance with increasing computational resources. Finally, we evaluated computational cost versus performance gain as EPO scales with agent numbers (Supplementary Material). Each aspect is discussed in detail.
@@ 5.1 Experimental Setup
Tasks We evaluate our algorithm across eight challenging environments covering a diverse range of tasks, including manipulation [34]—Single-Arm Regrasping (M:SG), Single-Arm Throwing (M:ST), Two-Arm Regrasping (M:TG), Two-Arm Reorientation (M:TR), and Two-Arm Reorientation with Obstacle (M:TO); locomotion—ANYmal Walking (L:A)[16] and Unitree A1 Parkour (L:U)[4]; and classic control—Humanoid from the DeepMind Control Suite (D)[45]. A visual overview of these tasks is shown in Figure2; additional details are provided in the Appendix.
Following prior works [34, 41], we use the number of successes per episode as the performance metric for manipulation tasks. For L:A and D, we report the cumulative rewards as defined in their original benchmarks. For L:U, we use the average terrain difficulty level achieved by all A1 agents during an episode, which serves as a proxy for success rate.
Figure 3: Performance curves of EPO compared to SAC, PQL, PPO, SAPG, PBT, and CEM-RL baselines. EPO demonstrates superior sample efficiency and higher asymptotic performance, particularly on challenging tasks.
Baselines We evaluate our approach against state-of-the-art RL methods tailored for large-scale training. The comparisons include off-policy algorithms—SAC [14] and Parallel Q-Learning [24]; on-policy methods—PPO [38]; hybrid-policy methods—SAPG [41]; and population-based EvoRL methods—PBT [34] and CEM-RL [35]. We use the original implementations of each method without introducing algorithm-level modifications. Please refer to the supplementary material for additional details.
For a fair comparison, all algorithms are trained using the same number of environments ($N=24{,}576$) and identical network architectures. For methods incorporating a population-based component (SAPG, PBT, CEM-RL), we standardize the population size to 64. Since the impact of the number of agents on final performance can vary, we also report results using each method’s original agent configuration for completeness (e.g., PBT: 8 agents; SAPG: 6 agents). Each experiment is run with 5 different random seeds, and we report the mean and standard error in the plots. Further ablation studies on the effect of population size are presented in Section 5.3.
Task SAC [14] PQL [24] PPO [38] SAPG[41] SAPG 6[41] PBT [34] PBT 8 [34] CEM-RL[41] EPO (Ours) M:SG $0.1\pm 0.00$ $3.1\pm 0.31$ $0.1\pm 0.11$ $22.4\pm 7.60$ $31.6\pm 2.07$ $6.8\pm 5.01$ $24.8\pm 1.17$ $1.7\pm 0.95$ 32.1$\pm$ 1.21 M:ST $0.0$ $2.5\pm 0.43$ $12.4\pm 3.11$ $10.8\pm 1.70$ $11.5\pm 4.15$ $2.2\pm 1.54$ $9.4\pm 2.73$ $0.0$ 16.4 $\pm$ 0.99 M:TG $0.0$ $1.7\pm 0.30$ $0.3\pm 0.15$ $1.4\pm 1.04$ $4.1\pm 2.04$ $3.1\pm 1.24$ $6.7\pm 1.43$ $0.0$ 24.1 $\pm$ 2.55 M:TR $0.0$ $0.5\pm 0.22$ $0.7\pm 0.45$ $4.4\pm 4.17$ $2.2\pm 1.17$ 0.0 $1.7\pm 1.46$ $0.0$ 35.8 $\pm$ 1.64 M:TO $0.0$ $0.0$ $0.5\pm 0.43$ $2.6\pm 0.76$ $2.4\pm 0.52$ $0.0$ $0.0$ $0.0$ 24.7 $\pm$ 7.06 L:A $64.2\pm 6.60$ $65.12\pm 3.2$ $67.7\pm 1.71$ $69.8\pm 1.81$ $68.9\pm 1.22$ $55.2\pm 1.72$ $64.7\pm 6.16$ $67.3\pm 2.01$ 74.0 $\pm$ 6.65 L:U $0.0$ $0.1\pm 0.03$ $0.6\pm 0.32$ $0.6\pm 0.45$ $0.5\pm 0.43$ $0.0$ $0.3\pm 0.31$ $0.1\pm 0.09$ 1.0 $\pm$ 0.45 D $5254.8\pm 133.95$ $8554.8\pm 593.32$ $9231.4\pm 88.47$ $11411.9\pm 790.88$ $11524.1\pm 540.23$ $7510.2\pm 117.86$ $9490.52\pm 41.09$ $7040.5\pm 103.09$ 11686.5 $\pm$ 773.7
@@
Figure 4: Ablation and scaling results for EPO. (a) Ablation of agent number ($K=8,16,32,64$) with a fixed total number of environments ($N=24{,}576$). As the number of agents increases, EPO’s performance improves. (b) Training and scaling curves with increasing environments. With a fixed number of environments per agent (384), total performance improves as environments scale.

### v3

@@ Abstract
Figure 1: Evolutionary Policy Optimization (EPO) integrates genetic algorithms with policy gradients. A population of agents, represented by latent genes and sharing one actor-critic network, interacts with the environment. A master agent is trained on aggregated experiences from all agents to improve efficiency and stability.
@@ 1 Introduction
Reinforcement learning (RL) has become a powerful paradigm for training autonomous decision-making agents across a wide range of domains, including games Silver et al. (2016); Berner et al. (2019); Vinyals et al. (2019); Mnih (2016), robotics Kalashnikov et al. (2018); Akkaya et al. (2019); Kumar et al. (2021); Hwangbo et al. (2019), and large-language-model alignment Ouyang et al. (2022); Guo et al. (2025); Team et al. (2023). Like all trial-and-error methods, RL is inherently sample-inefficient, and the problem is particularly acute for on-policy algorithms, which discard past experience after each update.
@@
Despite this drawback, on-policy model-free methods—commonly referred to as policy gradients Schulman et al. (2017); Mnih (2016)—are widely adopted in real-world applications Silver et al. (2016); Berner et al. (2019); Kalashnikov et al. (2018); Akkaya et al. (2019); Ouyang et al. (2022). Their popularity stems from their ability to achieve higher asymptotic returns and stable performance, making them the method of choice in data-rich settings such as simulations and game engines Makoviychuk et al. , where large batches are easy to generate. Unfortunately, recent works Singla et al. (2024); Petrenko et al. (2023) show that policy-gradient algorithms do not scale well with larger batch sizes: because data are collected from the current policy, adding more parallel environments does not guarantee greater diversity. Instead, the data distribution quickly converges, so simply running more environments on additional GPUs yields highly correlated trajectories rather than useful variety.
@@
Evolutionary algorithms (EAs) contribute the complementary strength of explicit diversity through random perturbations, making them natural partners for RL. Their combination—evolutionary RL (EvoRL) Moriarty et al. (1999); Salimans et al. (2017)—maintains a population of policies and refines them with evolutionary operators. Yet EvoRL has not been widely adopted. Classic EvoRL relies on gradient-free evolutionary strategies (ES) Khadka & Tumer (2018); Chen et al. (2019); Zheng & Cheng (2023), whose simplest variants reduce to random search Mania et al. (2018): they parallelize well but remain extremely sample-inefficient. Methods such as CEM-RL Pourchot & Sigaud (2019) introduce off-policy updates to improve efficiency but degrade asymptotic performance owing to distribution shift. Population-Based Training (PBT) Petrenko et al. (2023); Jaderberg et al. (2017) ensembles multiple on-policy agents, yet each policy evolves in isolation, wasting the opportunity to amortize experience across the population.
@@
Can we design an algorithm that (i) scales with the number of parallel environments, (ii) learns efficiently from shared experience, and (iii) preserves the high asymptotic returns and stability of on-policy methods? We answer in the affirmative. We introduce Evolutionary Policy Optimization (EPO), a novel policy-gradient algorithm that merges the strengths of EA and policy gradients: EA supplies diverse experience during massively parallel training, and experience from all agents is amortized to train a master agent via the split-and-aggregate policy-gradient (SAPG) formulation Singla et al. (2024), which performs off-policy updates with importance sampling to fold follower data into the master’s on-policy set.
@@
EPO maintains a population of agents alongside a master agent, all sharing the same network weights (analogous to gene-expression rules). This design prevents parameter bloat as the population scales, enabling efficient training and storage of large networks. Each agent is conditioned on a unique latent variable (“gene”), ensuring behavioral diversity while preserving coherence. Periodically, Darwinian selection Darwin (2023) is applied: low performers are removed, and elite agents are retained. These elites undergo crossover and mutation, injecting controlled diversity without excessive divergence. All agents are updated synchronously with policy gradients using shared reward signals. In every iteration, EPO aggregates experience from all agents into the master via SAPG updates Singla et al. (2024), enabling the master to learn from diverse, high-quality data and fully exploit massive parallelism.
@@
We evaluate EPO on several challenging and realistic environments spanning manipulation (Isaac Gym Allegro-Kuka Makoviychuk et al. ), locomotion (ANYmal Walking Hutter et al. (2016), Unitree A1 Parkour Cheng et al. (2024); Zhuang et al. (2023)), and classic control benchmarks (DeepMind Control Suite Tassa et al. (2018)). Across this suite, EPO significantly outperforms prior state-of-the-art RL methods. Moreover, our scaling-law analysis shows that EPO effectively scales with increased computational resources, making it a promising approach for large-scale RL training.
@@ 2 Related Work
On-policy approaches Sutton et al. (1999) update the policy using data collected by the current policy, ensuring that the training data remains closely aligned with the policy’s behavior. This strategy often leads to higher asymptotic rewards and stable performance, but at the cost of sample inefficiency, as past experiences from older policies are discarded. Despite this drawback, on-policy methods are the preferred choice in settings where data is abundant, such as simulations and game engines, where large training batches can be efficiently generated. In practice, on-policy RL has proven highly effective for training autonomous decision-making agents in games Silver et al. (2016); Berner et al. (2019); Vinyals et al. (2019); Mnih (2016), robotics Kalashnikov et al. (2018); Akkaya et al. (2019); Kumar et al. (2021); Hwangbo et al. (2019), and more recently, in developing reasoning capabilities in large language models (LLMs) Ouyang et al. (2022); Guo et al. (2025); Team et al. (2023). Two widely used algorithms in this category are Trust Region Policy Optimization (TRPO) Schulman et al. (2015) and Proximal Policy Optimization (PPO) Schulman et al. (2017), both designed to stabilize learning by constraining the magnitude of policy updates.
@@
Unlike on-policy algorithms, off-policy methods allow for policy updates using data collected from different or older policies, enabling more efficient reuse of past experiences. Classic approaches such as Q-learning and its deep variant Deep Q-Network (DQN) Mnih (2013) rely on a replay buffer to store state-action transitions for more stable learning. More recent continuous control algorithms, such as Deep Deterministic Policy Gradient (DDPG) Lillicrap (2015); Silver et al. (2014), Twin Delayed DDPG (TD3) Fujimoto et al. (2018), and Soft Actor-Critic (SAC) Haarnoja et al. (2018), extend these ideas within an actor-critic framework, further improving stability and performance. While off-policy methods are typically more sample-efficient than on-policy methods, they often suffer from distribution shift and overestimation of action values, which can hinder stable learning. To effectively amortize experience across all policies in a population while ensuring stable updates, we adopt the Split and Aggregate Policy Gradient (SAPG) formulation Singla et al. (2024), which performs off-policy updates via importance sampling, enabling the aggregation of experience from individual follower policies into the on-policy data of a master policy.
@@
Another related line of work is Evolutionary Reinforcement Learning (EvoRL) Moriarty et al. (1999), which integrates population-based search with policy optimization to explore a broad range of candidate solutions. Historically, evolutionary computation (EC) has been employed to optimize neural network weights Tang et al. (2020); Choromanski et al. (2018; 2019), architectures Gaier & Ha (2019), and hyperparameters Jaderberg et al. (2017); Snoek et al. (2012); Vincent et al. (2024). For instance, OpenAI ES Chrabaszcz et al. (2018) applied a streamlined evolution strategy to Atari games. More recently, researchers have integrated gradient-based updates alongside evolutionary operations (e.g., mutation and crossover), improving scalability and leveraging the representational power of deep neural networks Pourchot & Sigaud (2019); Khadka et al. (2019). For example, Khadka & Tumer (2018) extends Deep Deterministic Policy Gradient (DDPG) by introducing evolutionary operations on a population of policies, combining the exploratory benefits of evolution with gradient-based optimization for more robust learning. However, these methods remain vulnerable to distribution shift, which undermines their asymptotic performance. In this work, we propose to combine the strengths of evolutionary search with on-policy updates, enabling higher asymptotic rewards while maintaining high sample efficiency.
@@ 3 Preliminaries
is one of the most well-known Evolutionary Algorithms (EAs). It relies on three key operators: selection, mutation, and crossover. A core concept in genetic algorithms is the population, where each individual represents a potential solution evaluated using a fitness function. The classic genetic algorithm emphasizes crossover as the primary mechanism for exploration Hoffmeister & Bäck (1990) and is usually applied to deal with the problem of hyper-parameter tuning in reinforcement learning.
@@ 4 Approach
To further enhance learning efficiency, we adopted a hybrid policy update strategy Singla et al. (2024). The master agent updates its parameters using both (i) on-policy data collected in the most recent episode and (ii) off-policy data sampled from the trajectories of the entire population. Meanwhile, each non-master agent is updated using only its own on-policy data. This hybrid learning process ensures both efficient exploration and sample reuse, making EPO well-suited for large-scale reinforcement learning. In the following sections, we describe each component of EPO in detail.
@@ 4.1 Genetic Algorithm for Agent Selection
The design philosophy of our algorithm is twofold: (1) to generate diverse yet bounded data for training the master agent and (2) to ensure that non-master agents are updated efficiently, enabling them to produce useful yet diverse data. Drawing inspiration from biological evolution, we propose initializing a population of agents ($\{\pi_{k}\}_{k=1}^{K}$), where each agent is associated with a unique latent embedding ($\{\phi_{k}\in\mathcal{R}^{N}\}_{k=1}^{K}$). All agents share a common actor-critic network, parameterized by $\theta$ and $\psi$, respectively. The actor is conditioned on both the current state and its latent embedding to generate actions, while the value function is similarly conditioned to estimate value functions. Prior works have also leveraged latent-conditioned policies to encourage diverse behaviors Ebert et al. (2021); Eysenbach et al. (2018). Among these agents, one is designated as the master agent—for simplicity, we define the master agent as $\pi_{1}(\theta,\psi,\phi_{1})$, though the general framework allows for alternative assignments.
Figure 2: Evolutionary Policy Optimization
2: Initialize $K$ empty cyclic replay buffers $\{S\}_{i=1}^{K}$
3: Initialize NN environments $\{E\}_{i=1}^{N}$
4:for k = 1, …, $K$ do
5:  $E^{k}=\{E\}_{i=\frac{(k-1)N}{K}}^{\frac{kN}{K}}$ $\triangleright$ Assign environments
6:end for
7:for iteration=1,2,… do
8:  for k = 2, …, $K$ do
9:   $f_{k}\leftarrow Evaluate(E^{k},\pi_{k})$
10:  end for
11:  if Eq 3 then
12:   Rank population by fitness scores $f_{k}$
13:   Select top-$x$ as elites $X$, append to $Y$
14:   while |Y|<K1|Y|<K-1 do
[…]
17:    Append $\pi(\theta,\psi,\phi^{\prime\prime})$ to $Y$
18:   end while
19:   $\{\pi_{i}\}_{i=2}^{K}\leftarrow Y$
20:  end if
21:  for k = 1, …, $K$ do
[…]
23:  end for
24:  Sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ to get $S_{1}^{\prime}$
25:  $L\leftarrow$ OffPolicyLoss($S_{1}^{\prime}$)
26:  for k = 1, …, $K$ do
27:   $L\leftarrow L+$ OnPolicyLoss($S_{k}$)
28:  end for
29:  Update $\theta,\psi,\{\phi_{k}\}^{K}_{k=1}$
30:end for
31: Return $\theta,\psi,\phi_{1}$
@@ 4.2 Hybrid-Policy Optimization
Different agents generate diverse data through varied behaviors, which can be beneficial if the master agent can effectively learn from all available data. However, since data distributions collected by different agents may vary significantly, incorporating off-policy data requires careful correction. Following Meng et al. (2023); Singla et al. (2024), we apply importance sampling to reweight updates from different policies. Specifically, we sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ collected by agents $\{\pi_{k}\}_{k=2}^{K}$ to get $S^{\prime}_{1}$. To update the master agent $\pi_{1}$ using $S^{\prime}_{1}$, we define the off-policy loss as:
@@ 5 Experiments
Figure 3: We evaluate our algorithm on eight challenging environments that span a diverse set of tasks, including manipulation Petrenko et al. (2023), locomotion Cheng et al. (2024), and classic control benchmarks Tassa et al. (2018)
We evaluate our method on eight challenging and realistic environments, including manipulation Petrenko et al. (2023), locomotion Hutter et al. (2016); Cheng et al. (2024), and classic control benchmarks Tassa et al. (2018), comparing it against state-of-the-art (SOTA) approaches. Additionally, we conduct ablation studies to assess the contribution of each component. To demonstrate scalability, we analyze our method’s performance with increasing computational resources. Finally, we evaluated computational cost versus performance gain as EPO scales with agent numbers (Supplementary Material). Each aspect is discussed in detail.
@@ 5.1 Experimental Setup
Tasks We evaluate our algorithm across eight challenging environments covering a diverse range of tasks, including manipulation Petrenko et al. (2023)—Single-Arm Regrasping (M:SG), Single-Arm Throwing (M:ST), Two-Arm Regrasping (M:TG), Two-Arm Reorientation (M:TR), and Two-Arm Reorientation with Obstacle (M:TO); locomotion—ANYmal Walking (L:A)Hutter et al. (2016) and Unitree A1 Parkour (L:U)Cheng et al. (2024); and classic control—Humanoid from the DeepMind Control Suite (D)Tassa et al. (2018). A visual overview of these tasks is shown in Figure3; additional details are provided in the Appendix.
Following prior works Petrenko et al. (2023); Singla et al. (2024), we use the number of successes per episode as the performance metric for manipulation tasks. For L:A and D, we report the cumulative rewards as defined in their original benchmarks. For L:U, we use the average terrain difficulty level achieved by all A1 agents during an episode, which serves as a proxy for success rate.
Figure 4: Performance curves of EPO compared to SAC, PQL, PPO, SAPG 64, PBT 8, and CEM-RL baselines. We plot best-performing agent numbers for SAPG (64) and PBT (8). EPO shows superior sample efficiency and higher asymptotic performance, particularly on challenging tasks.
Baselines We evaluate our approach against state-of-the-art RL methods tailored for large-scale training. The comparisons include off-policy algorithms—SAC Haarnoja et al. (2018) and Parallel Q-Learning Li et al. (2023); on-policy methods—PPO Schulman et al. (2017); hybrid-policy methods—SAPG Singla et al. (2024); and population-based EvoRL methods—PBT Petrenko et al. (2023) and CEM-RL Pourchot & Sigaud (2019). We use the original implementations of each method without introducing algorithm-level modifications. Please refer to the supplementary material for additional details.
For a fair comparison, all algorithms are trained using the same number of environments (N=24,576N=24{,}576) and identical network architectures. For methods incorporating a population-based component (SAPG, PBT, CEM-RL), we standardize the population size to 64. Since the impact of the number of agents on final performance can vary, we also report results using each method’s original agent configuration for completeness (e.g., PBT: 8 agents; SAPG: 6 agents) in the Table 1. Each experiment is run with 5 different random seeds, and we report the mean and standard error in the plots. Further ablation studies on the effect of population size are presented in Section 5.3.
[…]
@@
Figure 5: Ablation and scaling results for EPO. (a) Ablation of agent number ($K=8,16,32,64$) with a fixed total number of environments (N=24,576N=24{,}576). As the number of agents increases, EPO’s performance improves. (b) Training and scaling curves with increasing environments. With a fixed number of environments per agent (384), total performance improves as environments scale.